he plant started to purify its wastewater before discharging it into the lake. How long will it take for the concentration of the chemical to reduce to 1% of its steady-state value, if its half-life is 2 years?
Answer #1

The formula is given as A= A0(1/2)^(t/h)

h is half life and A is the final concentration.

A= 1, A0 = 100

1 = 100 (0.5)^(t/2)

0.01 = 0.5^(t/2)

Taking ln on both sides,

ln (0.01)= (t/2) ln(0.5)

-4.605 = (t/2) (-0.693)

t = 13.29 hrs

Time taken for the decay = 13.29 hrs